OpenStudy (anonymous):

The formula for the volume of a sphere is V= 4/3 pie r^3. Find the radius, to the nearest hundredth, of a sphere with a volume of 15 in.3.

OpenStudy (anonymous):

This is what the question looks like

OpenStudy (campbell_st):

ok... so the task is to make r the subject of the equation \[15 = \frac{4}{3} \pi r^3\] multiply both sides by 3 then divide both sides by 4 gives \[\frac{45}{4} = \pi r^3\] divide both sides by pi \[\frac{45}{4\pi} = r^3\] take the cube root to find the answer